OpenStudy (hulagirlmb):

Algebra 1 help: Write the equation in standard form. 8-4(x-y)=2x+6

ganeshie8 (ganeshie8):

standard form looks like below : \[\rm Ax + By = C\]

OpenStudy (hulagirlmb):

I think that it's 6x+4y=2

ganeshie8 (ganeshie8):

very close, but nope.

OpenStudy (hulagirlmb):

6x-4y=2 maybe?

ganeshie8 (ganeshie8):

thats right but i feel you're just guessing hmm

OpenStudy (hulagirlmb):

That was kinda guessing. I knew it was one of those 2 answers

ganeshie8 (ganeshie8):

rearranging it in standard form is pretty easy, you just need to know how to distribute and combine like terms

OpenStudy (hulagirlmb):

Ok thank you

ganeshie8 (ganeshie8):

\[\rm 8-4(x-y)=2x+6\] The first thing you want to do is to distribute that -4 into the parenthesis, you get : \[\rm 8-4x+4y=2x+6\]

OpenStudy (hulagirlmb):

Yeah I did that

ganeshie8 (ganeshie8):

You're ready to combine like terms now, just bring all variables to one side and constants to other side

OpenStudy (hulagirlmb):